# Dynamical Dark Energy from Lattice Quantum Gravity

### v1

@@ IV.6 Implications for the running vacuum model
leads to an equation of state $w$ that ensures this is the case. The good agreement between our two methods of obtaining $\nu$ (where $\nu=\nu^{\prime}$ when $\tilde{\nu}=0$), provides very strong evidence that Eq. (75) is the correct form for the vacuum pressure that emerges from the lattice. This has important implications when applying this model to cosmology, as discussed in the next section.
@@ V Making contact with Cosmology
An important feature of this model is that the vacuum pressure given by Eq. (75) ensures that the vacuum energy is covariantly conserved independently of the matter sector. If this were not the case and the decrease in vacuum energy was accompanied by a transfer of energy into matter or radiation, this would lead to strong violations of experimental constraints. For example, if the vacuum decayed to radiation, this would lead to a large deviation of the cosmic microwave background from the $\Lambda$CDM expectation and would not be consistent with observations. If the vacuum decayed to cold dark matter during the radiation era, then far too much dark matter would be produced to be compatible with the current energy budget of the universe. That the vacuum energy is separately covariantly conserved is a highly-nontrivial result emerging from our lattice calculations, and it ensures that our result does not significantly modify the successes of the standard $\Lambda$CDM picture. It is worth emphasizing that there is no need to put this result in by hand.
@@
where we have assumed that the radiation density can be ignored in the present era. This expression can be rewritten as [67]
@@ VI Conclusion and Outlook
New ensembles and methods for determining the lattice spacing enable us to test the de Sitter nature of our geometries at a level of precision not previously possible, and a careful study shows a deviation from this picture. Non-trivial vacuum dynamics is seen, with a power law running of the cosmological constant. Good agreement across four lattice spacings suggests that the running is not a lattice artifact. Consistency checks also provide strong evidence that the running seen in the simulations is not an artifact, foremost among these the consistency of the lattice results with covariant energy conservation. The running vacuum model that is singled out by our lattice calculation is summarized by Eqs. (73), (74), and (75). This is a one-parameter model, since $\Lambda_{0}$ must be taken from experiment, while $\nu$ and its running are fixed from first principles by the analysis presented here. An additional parameter, $\tilde{\nu}$, appearing in a more general version of the running vacuum model, is consistent with zero and is constrained to be much smaller than $\nu$ by our analysis.
@@
The lattice calculation presented here could be improved in a number of ways. Going to larger, finer lattices would give better control over discretization effects, especially the trending of $\nu$ towards negative values if the volume is taken too large at too coarse of a lattice spacing, as seen in Figure 16. A better determination of the relative lattice spacings would also help, and additional runs to help pin these down are in progress. Unquenching the calculation is also a high priority for future work. Including the quantum fluctuations of light fields, for example those of the Standard Model, would lead to a back-reaction on the geometry, which could potentially alter $\nu$. We expect that this would change $\nu$ by an ${\cal O}(1)$ constant, not by orders of magnitude, but it is important to do the simulations in order to test this assumption. According to a semiclassical calculation [28], heavy particles would contribute directly to $\nu$ in proportion to their mass squared. Thus, the existence of a large number of GUT-scale particles could potentially modify $\nu$ from our result, which ought to be regarded as the pure gravity prediction. Unquenching could also impact the predictions of the pure-gravity result in a more dramatic way if threshold effects are present, as discussed below.
@@ Acknowledgements.
The authors thank Simon Catterall and Scott Watson for valuable discussions and for comments on the manuscript. MD and JL were supported by the U.S. Department of Energy (DOE), Office of Science, Office of High Energy Physics under Award Number DE-SC0009998. JUY was supported by the U.S. Department of Energy grant DE-SC0019139, and by Fermi Research Alliance, LLC under Contract No. DE-AC02- 07CH11359 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics, and by the Department of Energy through the Fermilab Theory QuantiSED program in the area of “Intersections of QIS and Theoretical Particle Physics.” MS acknowledges support by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. Computations were performed in part on the Syracuse University HTC Campus Grid and were supported by NSF award ACI-1341006. This research was enabled in part by support provided by ACENET, Calcul Québec, Compute Ontario, and the Digital Research Alliance of Canada. The authors acknowledge support by the state of Baden-Württemberg through bwHPC.

### v2

@@ IV.6 Implications for the running vacuum model
leads to an equation of state $w$ that ensures this is the case. The absence of an emergent matter sector in our simulations suggests that the vacuum sector independently maintains covariant energy conservation, and that the vacuum pressure adjusts itself so that this is the case. This has important implications when applying this model to cosmology, as discussed in the next section.
@@ V Making contact with Cosmology
An important feature of this model is that the vacuum pressure given by Eq. (75) ensures that the vacuum energy is covariantly conserved independently of the matter sector. If this were not the case and the decrease in vacuum energy was accompanied by a transfer of energy into matter or radiation, this would lead to strong violations of experimental constraints. For example, if the vacuum decayed to radiation, this would lead to a large deviation of the cosmic microwave background from the $\Lambda$CDM expectation and would not be consistent with observations. If the vacuum decayed to cold dark matter during the radiation era, then far too much dark matter would be produced to be compatible with the current energy budget of the universe. Thus, the separate covariant energy conservation of the vacuum sector ensures that the vacuum dynamics does not significantly modify the successes of the standard $\Lambda$CDM picture.
@@
where we have assumed that the radiation density can be ignored in the present era. This expression can be rewritten as [29, 69]
@@ VI Conclusion and Outlook
New ensembles and methods for determining the lattice spacing enable us to test the de Sitter nature of our geometries at a level of precision not previously possible, and a careful study shows a deviation from this picture. Non-trivial vacuum dynamics is seen, with a power law running of the cosmological constant. Good agreement across four lattice spacings suggests that the running is not a lattice artifact. Consistency checks also provide strong evidence that the running seen in the simulations is not an artifact. The running vacuum model that is singled out by our lattice calculation is summarized by Eqs. (73), (74), and (75). This is a one-parameter model, since $\Lambda_{0}$ must be taken from experiment, while $\nu$ and its running are fixed from first principles by the analysis presented here. An additional parameter, $\tilde{\nu}$, appearing in a more general version of the running vacuum model, is consistent with zero and is constrained to be much smaller than $\nu$ by our analysis.
@@
The lattice calculation presented here could be improved in a number of ways. Going to larger, finer lattices would give better control over discretization effects, especially the trending of $\nu$ towards negative values if the volume is taken too large at too coarse of a lattice spacing, as seen in Figure 16. A better determination of the relative lattice spacings would also help, and additional runs to help pin these down are in progress. Unquenching the calculation is also a high priority for future work. Including the quantum fluctuations of light fields, for example those of the Standard Model, would lead to a back-reaction on the geometry, which could potentially alter $\nu$. We expect that this would change $\nu$ by an ${\cal O}(1)$ constant, not by orders of magnitude, but it is important to do the simulations in order to test this assumption. According to a semiclassical calculation [29, 30], heavy particles would contribute directly to $\nu$ in proportion to their mass squared. Thus, the existence of a large number of GUT-scale particles could potentially modify $\nu$ from our result, which ought to be regarded as the pure gravity prediction. Unquenching could also impact the predictions of the pure-gravity result in a more dramatic way if threshold effects are present, as discussed below.
@@ Acknowledgements.
The authors thank Claude Bernard, Simon Catterall, and Scott Watson for valuable discussions and for comments on the manuscript. MD and JL were supported by the U.S. Department of Energy (DOE), Office of Science, Office of High Energy Physics under Award Number DE-SC0009998. JUY was supported by the U.S. Department of Energy grant DE-SC0019139, and by Fermi Research Alliance, LLC under Contract No. DE-AC02- 07CH11359 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics, and by the Department of Energy through the Fermilab Theory QuantiSED program in the area of “Intersections of QIS and Theoretical Particle Physics.” MS acknowledges support by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. Computations were performed in part on the Syracuse University HTC Campus Grid and were supported by NSF award ACI-1341006. This research was enabled in part by support provided by ACENET, Calcul Québec, Compute Ontario, and the Digital Research Alliance of Canada. The authors acknowledge support by the state of Baden-Württemberg through bwHPC.